Mr. Baker is mixing a 16% saline solution with a 21% saline solution. He would like to obtain 25 gram mixture that is 20% saline. How many grams of each will be needed?
Answer by Alan3354(69443)

Mr. Baker is mixing a 16% saline solution with a 21% saline solution. He would like to obtain 25 gram mixture that is 20% saline. How many grams of each will be needed?
----------
s = amount of 16%
t = amount of 21%
---
s + t = 25 (total solution)
16s + 21t = 25*20 (total salt)